lexxii:

I honestly think I'm just so overly tired but I'm not getting this. Could someone help me out?

lexxii:

@Shadow if you have the time would you mind walking me through this?

lexxii:

@JustSaiyan

Hero:

@lexxii are you still working on this?

lexxii:

I did some work and submitted it in but I don't believe I did it right at all. So I was hoping someone would still explain what I would need to do and walk me through it.

Hero:

Sure thing.

Hero:

@lexxii

Hero:

So they have given us the Demand Function $$D(x) = 100000\sqrt{2x}$$ and they want us to find the inverse of \(D(x)\) which will give us \(S(x)\) the Supply Function. @lexxii do you remember how to find the inverse of a function?

lexxii:

I don't

Hero:

Okay, so to take the inverse of a function, we usually do is do the following: 1. Replace \(D(x)\) with \(y\). 2. Swap variables \(x\) and \(y\). 3. Then manipulate the equation until we have \(y\) in terms of \(x\). For example, suppose we need to find the inverse of $$f(x) = 3(x + 2)$$ 1. \(y = 3(x + 2)\) 2. \(x = 3(y + 2)\) 3. Divide both sides by 3 then subtract 2 from both sides to get the inverse. $$\dfrac{x}{3} - 2 = y$$

Hero:

@lexxii would you like to try taking the inverse of \(D(x)\)?

lexxii:

So is all I have to do is find the inverse? That is what I was doing but I wasn't sure if I was doing it right but from your example I was. So after I find the inverse is that all I would need to do?

Hero:

It's not "all" that you need to do, but it's the first step.

lexxii:

So after I find the inverse what do I need to do? I'm fine with doing all the work and what not but what are all the steps I need to follow?

Hero:

Well it does say to also sketch and describe your product. Think of something that you like that you would love to sell. Then make sure whatever you're describing fits within the parameters of the supply and demand functions which basically would involve however much of the product you are able to manufacture over a one year period.

lexxii:

Yes that was one of the things I was able to knock out pretty quickly.

Hero:

Did you find the inverse of \(D(x)\) yet?

lexxii:

I was able to get most of it done. We got a message from our teacher saying she wanted to go over this with us tomorrow and she will take us step by step through it. Thank you for the help as always Hero:)

Hero:

You're welcome